# Future proofing network encryption technology (and securing critical infrastructure data) with continuous-variable quantum key distribution

### v1

@@ IV.2 Practical and real-time data processing
As presented in the previous section, the bottlenecks in the implementation of the CVQKD protocol stack primarily come from the DSP at RX as well as the IR and PA procedures. Also, for a fully autonomous operation, the data processing pipelines at both TX and RX require synchronization.
@@
The main challenge that then remains to solve is how to pick the appropriate input parameters such as the code rate, the log likelihood ratio factor, etc. that ensure a single-shot reconciliation with the simultaneous assurance of a high $\beta$ at a reasonable FER. The overall IR performance is fairly sensitive to the choice of these parameters, which also need to be frequently adjusted in case of channels with a fluctuating transmittance. For this, we are currently working on an algorithm that uses the ancillary signals to enable a dynamic calculation of these input parameters which can then make the IR part of the classical data processing stack meet all the aforementioned requirements. This would allow developing a feedback between TX and RX to optimize the modulation strength, thus maximising the SKR for the given channel parameters.
V Conclusion & Outlook
In conclusion, we have reported the field deployment of a semi-autonomous continuous variable quantum key distribution (CVQKD) system that powers a network encryption scheme to establish quantum-secure data transfer links at two locations in Denmark. More specifically, system operation in a power substation (critical infrastructure) environment and the ability to perform quantum communication over different types of channels that are a combination of underground and aerial fibers is showcased. The CVQKD system operates at 1550 nm while the network encryptors operate at 1300 nm and use the same (wavelength-multiplexed) channel for the transport of both quantum states as well as encrypted data. Despite the several orders of magnitude difference in optical powers, the scheme is able to have a positive secret key yield, exhibiting another successful quantum-classical integration example. Such a crypto-agile platform can be used to future proof the security of network data infrastructure, such as data center interconnections.
VI Acknowledgements

### v2

@@ 4.2 Practical and real-time data processing
As presented in the previous section, the bottlenecks in a real time implementation of the CVQKD protocol stack primarily come from the DSP at RX as well as the IR and PA procedures. Also, for a fully autonomous operation, the data processing pipelines at both TX and RX require synchronization.
@@
The main challenge that then remains to solve is how to pick the appropriate input parameters such as the code rate, the log likelihood ratio factor, etc. that ensure a single-shot reconciliation with the simultaneous assurance of a high $\beta$ at a reasonable FER. The overall IR performance is fairly sensitive to the choice of these parameters, which also need to be frequently adjusted in case of channels with a fluctuating transmittance [34]. For this, we are currently working on an algorithm that uses the ancillary signals to enable a dynamic calculation of these input parameters which can then make the IR part of the classical data processing stack meet all the aforementioned requirements. This would additionally facilitate a feedback between TX and RX to optimize the modulation strength, thus maximising the SKR for the given channel parameters.
5 Conclusion & Outlook
In conclusion, we have reported field trials of a semi-autonomous continuous variable quantum key distribution (CVQKD) system that powers a network encryption scheme to establish quantum-secure data transfer links at two locations in Denmark. More specifically, we successfully operated the CVQKD system outside the lab, with the devices located in a high electromagnetic noise environment of a power substation and performing quantum communication over combinations of underground and aerial fibers as the quantum channels. The CVQKD system operated at 1550 nm while the network encryptors operated at 1300 nm and used the same (wavelength-multiplexed) channel for the transport of both quantum states as well as encrypted data. Despite the several orders of magnitude difference in optical powers, the scheme was able to have a positive secret key yield on 3 of the 4 investigated channels, exhibiting another successful quantum-classical integration example. This showcases the technological readiness level of QKD for protecting critical infrastructure data in a reliable and robust way. Furthermore, such a quantum-classical platform can be used to future proof the security of network data infrastructure, such as data center interconnections.
Acknowledgements
@@
Disclosures
The authors declare no conflicts of interest.
Data availability
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.